If an employee deposits Rs. 2,000 at the end of each year into his company’s plan which pays 7% interest compounded quarterly, how much will he have in the account at the end of 5 years?
1
Expert's answer
2019-05-20T13:41:20-0400

Future value for annuity:

"FV=P\frac{(1+\frac{r}{n})^{nt}-1}{\frac{r}{n}}=2000\frac{(1+\frac{0.07}{4})^{20}-1}{\frac{0.07}{4}}=47403.22"